OpenStudy (anonymous):

An achitect is allowed 15 square yards of floor space to add a small bedroom to a house. Because of the room design in relationship to the existing structure, the width of the rectangular floor must be 7 yards less than two times the length. Find he length and width of the rectangular floor that the architect is permitted.

OpenStudy (anonymous):

we just solved it

OpenStudy (anonymous):

if l is the length, then the width is 2l-7 the area is 15 so l(2l-7)=15

OpenStudy (anonymous):

huh? no. the answer is l =5yards and width = 3 yards

OpenStudy (anonymous):

solve the quadratic equation I posted.

OpenStudy (anonymous):

my bad. didn't subtract the 15

OpenStudy (anonymous):

\[2l ^{2}-7l-15=0\]\[(2l+3)(l-5)=0\]\[\therefore l=5,(-3/2)\](-3/2) is extraneous, so \[length=5\] and \[width =5\times2-7=3\]